OpenStudy (anonymous):

Which expression is equivalent to 46 ⋅ 4−8? 1 over 4 to the 2nd power 1 over 4 to the 14th power 4^2 4^14

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

when we multiply same bases we should `ADD` their exponents \[\huge\rm x^m \times x^n=x^{m+n}\] and please use `^` for exponents

OpenStudy (anonymous):

ok

Nnesha (nnesha):

expression is \[\huge\rm 4^6 \times 4^{-8}\] bases are same so add their exponents

OpenStudy (anonymous):

-2?

Nnesha (nnesha):

yes right now flip the fraction to make it posisitve exponnet \[\huge\rm x^{-m} =\frac{ 1 }{ x^m }\]

Nnesha (nnesha):

ugh positive ** exponent **

OpenStudy (anonymous):

wait isnt it flipped already because the exponents already positive?

Nnesha (nnesha):

no it's not you just said it's -2

Nnesha (nnesha):

\[\huge\rm 4^{6+(-8)}= 4^{-2}\]

OpenStudy (anonymous):

hmm 2^4?

OpenStudy (welshfella):

lol Nnesha you are making money!!

OpenStudy (anonymous):

sorry im not the best with fracs and exponents

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ x^{-m} }{ 1}=\frac{ 1 }{ x^{m} }\] `flip the fraction` and change the sign that's it :D

Nnesha (nnesha):

lol welshfella

OpenStudy (anonymous):

ok zo the fractions was

OpenStudy (anonymous):

so*

Nnesha (nnesha):

it's \[4^{-2}~is ~ same ~ as ~\frac{ 4^{-2} }{ 1}\]

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

1/4^2

Nnesha (nnesha):

perfect!

OpenStudy (anonymous):

:D

Nnesha (nnesha):

D:

OpenStudy (anonymous):

so a?

Nnesha (nnesha):

no its D

OpenStudy (anonymous):

thats 4^14

Nnesha (nnesha):

so why did you ask ? lol ofc its a

OpenStudy (anonymous):

oh xDD

OpenStudy (anonymous):

Thanks ur awesome